# Large Language Models for Cryptocurrency Transaction Analysis: A Bitcoin Case Study

### v2

@@
Abstract.
Cryptocurrencies are widely used, yet current methods for analyzing transactions heavily rely on opaque, black-box models. These lack interpretability and adaptability, failing to effectively capture behavioral patterns. Many researchers, including us, believe that Large Language Models (LLMs) could bridge this gap due to their robust reasoning abilities for complex tasks.
In this paper, we test this hypothesis by applying LLMs to real-world cryptocurrency transaction graphs, specifically within the Bitcoin network. We introduce a three-tiered framework to assess LLM capabilities: foundational metrics, characteristic overview, and contextual interpretation. This includes a new, human-readable graph representation format, LLM4TG, and a connectivity-enhanced sampling algorithm, CETraS, which simplifies larger transaction graphs.
Experimental results show that LLMs excel at foundational metrics and offer detailed characteristic overviews. Their effectiveness in contextual interpretation suggests they can provide useful explanations of transaction behaviors, even with limited labeled data.
LLM, Cryptocurrency, Transaction Graph, Bitcoin Network
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Applied computing Network forensicsccs: Computing methodologies Machine learning
1. Introduction
@@
Large language models (LLMs) (chang2024survey, ) have significantly boosted the productivity of daily life and have a huge impact on the research community. LLMs broadened the boundaries of numerous fields, including natural language processing (NLP) (min2023recent, ; tinn2023fine, ), computer vision (CV) (wang2024visionllm, ; zhu2023minigpt, ), and application research (he2023large, ; 287298, ; singhal2023large, ). Moreover, the applications of LLMs extend beyond traditional domains, influencing areas with social and economic implications.
@@
One such area is the cryptocurrency ecosystem. Its growing adoption in finance, retail, and entertainment has led to a surge in transaction volumes. However, the expansion also exposes the ecosystem to risks, such as scams and money laundering, enabled by its decentralized and pseudoanonymous nature. Current analysis methods rely on black-box models and struggle with interpretability and adaptability. In this context, applying LLMs to analyze cryptocurrency transactions offers a promising approach to bridging these gaps. By leveraging their capacity to interpret complex patterns and behaviors, LLMs can help identify illicit activities and enhance cybercrime detection efforts.
@@
Although LLMs trained on massive datasets excel in NLP tasks, their application to graph analysis presents challenges due to structural differences between graph and text data. Recent studies investigated the possibilities of LLMs for handling graph data-related tasks, concluding affirmatively that they are capable of completing specific tasks with acceptable performance on graphs such as small graphs, citation graphs, or knowledge graphs (KGs) (guo2023gpt4graph, ; hu2023beyond, ; wang2023can, ; zhang2023llm4dyg, ). Nevertheless, measuring LLMs’ capability to understand and analyze cryptocurrency transaction graphs remains impractical. They contain different information compared with the other graph types such as KGs. Taking the Bitcoin transaction network as an example, the node represents the Bitcoin address or transaction, while the edge indicates the token flows among the address nodes and the transaction nodes (10375557, ).
@@
In addition, due to the input token limit of LLMs, how to efficiently feed larger graph data into LLMs to gain more information for potentially improving the quality of generated answers to various questions relevant to Bitcoin transaction graphs (e.g., address type prediction) continues to be an open question. To bridge the gaps in applying LLMs to transaction graph analysis, we study Bitcoin networks and address three research questions:
  • $\bullet$
@@
Contribution. Our work, for the first time, evaluates LLMs’ capabilities in analyzing transaction graphs for real-world cryptocurrencies, especially Bitcoin. We make the following contributions:
  • $\bullet$
@@
    We present a layered framework (§3) with three levels of understanding for measuring LLMs’ ability to analyze transaction graphs in cryptocurrency networks.
  • $\bullet$
@@
    We propose a text-based graph representation format, denoted LLM4TG (§3.2). It reduces redundant data and provides a human-readable syntax that naturally supports processing by LLMs.
  • $\bullet$
@@
    We design a Connectivity-Enhanced Transaction Graph Sampling algorithm (CETraS) (\$3.3) for graph summarization, targeting cutting off the less important nodes in middle-size transaction graphs while enhancing the critical connections.
  • $\bullet$
    We applied our framework to evaluate LLMs’ capabilities in real-world cryptocurrency transaction analysis using two datasets (BASD, BABD) extracted from the Bitcoin ledger and three mainstream models (GPT-3.5, GPT-4, GPT-4o), providing both quantitative assessment and qualitative analysis of their performance.
2. Technical Warmups
@@
Transaction graphs for cryptocurrencies. Cryptocurrency transaction graphs (chen2020understanding, ; weber2019anti, ; guo2019graph, ; yu2023predicting, ) represent the flow of digital currency between entities on blockchain networks such as Bitcoin. Each node in the graph typically represents a transaction or wallet address, and edges represent the transfer of cryptocurrency between these nodes. These graphs are crucial for analyzing the behavior of users, identifying patterns such as fraud or money laundering, and understanding the overall dynamics of the cryptocurrency market.
@@
Large language models. LLMs (chang2024survey, ) like OpenAI’s GPT (Generative Pre-trained Transformer) (achiam2023gpt, ) series are advanced AI models trained on vast amounts of text data. They excel in generating coherent and contextually relevant text based on the input they receive and can perform a variety of tasks without specific task-oriented training. While primarily designed for natural language processing, LLMs can be adapted to non-textual tasks such as analyzing structured data, including graphs (jin2023large, ; li2023survey, ). By converting data into a format that mimics natural language or structured prompts, researchers can leverage LLMs’ powerful generative and interpretative capabilities to perform complex analyses like those required for understanding cryptocurrency transaction graphs.
Within this paper, we focus on three mainstream LLM models: GPT-3.5 handles general tasks with a limited token capacity of 16K and training data up to mid-2021. GPT-4 increases the token limit to 128K and improves its ability to tackle complex analyses and research. GPT-4o maintains the same token capacity as GPT-4 while significantly enhancing efficiency.
@@
Tokens in LLMs. The term token indicates the smallest unit of text processed by LLMs. It can vary from individual characters, such as letters and punctuation, to more complicated units, such as words or subwords. LLMs analyze and generate text by processing token sequences. Each LLM has a maximum token limit per input, which poses a significant challenge when working with extensive data sets, such as large transaction graphs.
The token limit in LLMs like GPT-3 or GPT-4 (achiam2023gpt, ) constrains data analysis in a single query, posing challenges especially in analyzing large and complex cryptocurrency transaction graphs. This limitation may result in a loss of context or missing critical data. To overcome this, researchers employ strategies such as data compression to simplify input data, selective sampling to focus on the most relevant sections of graphs, custom tokenization to optimize data representation within token limits, and iterative processing to break the analysis into manageable parts.
3. Methodology
3.1. Quick Overview
@@
We present our framework in Fig. 1. The first step is to construct the Bitcoin transaction graph using the historical on-chain data. The second step is to extract corresponding subgraphs (or the relevant subgraphs sampled by CETraS, §3.3) and graph features to Bitcoin addresses with labels. Finally, raw graphs (formatted as LLM4TG, §3.2) and graph features are input to LLMs with proper prompts according to the tasks on different understanding levels.
Figure 1. LLM Evaluation Framework for Bitcoin Transaction
3.2. LLM4TG
@@
We introduce LLM4TG, a new format designed to optimize the analysis of transaction graphs using LLMs. This format is text-based and human-readable, minimizing syntactic noise/redundancy while reducing token usage and preserving data integrity.
@@
We denote T as transaction and A as address. <NodeID> represents the node’s ID, <Number> represents an integer. <Float> represents a real number. The syntax is displayed as follows:
@@
This format provides three key advantages for representing and analyzing transaction graphs: 1) It organizes nodes into type-specific layers, closely mirroring the structure of original transaction graphs; 2) It efficiently utilizes the limited token budget of LLMs by allowing more data to be encoded; 3) It improves the interpretability of graph data for LLMs by organizing node attributes into closely associated key-value pairs.
To further demonstrate the effectiveness of LLM4TG, we compared its token consumption with other formats for the same graphs, as shown in Fig. 7. This comparison reveals that LLM4TG experiences a more gradual increase in token usage and consistently stays within the GPT-4/4o token limit across various graph sizes, making it a more efficient format, especially for larger graphs.
3.3. CETraS
@@
In CETraS, nodes with lower importance are prioritized for elimination.to generate a subset of the nodes being preserved. The size of this retained subset is determined by a parameter that is flexible for specific demands. To maintain connectivity, paths connecting retained nodes are also preserved. Unlike other state-of-the-art graph summarization methods (hajiabadi2021graph, ; kang2022personalized, ; lee2020ssumm, ; lai2023optimized, ) focusing on keeping accuracy for structure-relevant queries or computations on large-scale graphs (billion node-level), CETraS concentrates on accurately conveying transaction-relevant information to LLMs for mid-scale graphs that typically contain thousands of nodes.
@@ 3       $I_{node}\leftarrow\left\{v:I_{node}[v]\text{ for each }v\text{ in }V(G)\right\}$;
4       $P_{node}\leftarrow\frac{1}{I_{node}}$, set $n_{0}$ probability to 0;
5       Normalize $P_{node}$ so that the sum equals 1;
6       $G_{subset}\leftarrow$ Sample from $V(G)$ with $P_{node}$ until $N_{target}$ or fewer nodes are chosen;
@@ 7      $G_{sampled}\leftarrow$ Initialize an empty graph;
8       foreach node $n$in $G_{subset}$ do
@@ 11                   Add node $m$ to $G_{sampled}$;
14            Add and all edges along $p$ to $G_{sampled}$;
15
@@ 14            Add and all edges along $p$ to $G_{sampled}$;
4. Evaluation and Analysis
4.1. Preparation
Dataset. We use two datasets for experiments. Both datasets are constructed from the Bitcoin transaction graph $G_{e}$. The data spans a 22-month period from 12 July 2019 to 26 May 2021.
  • $\bullet$
    BASD (xiang2022leveraging, ) includes eight types of subgraphs. Each subgraph, with a maximum of five hops and 3,000 nodes, is generated starting from a labeled Bitcoin address (denoted as $n_{0}$);
  • $\bullet$
    BABD (10375557, ) contains labeled Bitcoin addresses, each associated with 148 engineered features.
LLM selection. We select three LLMs, GPT-3.5, GPT-4, and GPT-4o (i.e., gpt-3.5-turbo, gpt-4-turbo, and gpt-4o) due to their capacity to handle larger token inputs (e.g., only 8,192 for LLaMA-3) and their superior performance. Both LLMs are accessed through application programming interfaces (APIs). Users can send requests with concrete prompts and receive responses.
Prompt engineering. We employ the few-shot prompting (wei2022chain, ) to interact with LLMs, where only a small number of examples are provided to guide the model in performing tasks effectively.
@@ 4.1. Preparation
In our experiments, we format raw transaction graphs as LLM4TG. We use CETraS in level 2 and level 3 due to the token limit in few-shot prompts. We apply GPT-3.5 only in level 3 feature-based classification, given its particularly tight token limits (Fig.7). Experiments are conducted on BASD dataset subsets, including data from its corresponding BABD addresses.
4.2. Level 1 - Foundational Metrics
We randomly select 50 transaction graphs, each with 10 chosen nodes (at least one per layer) for experiments. Besides, we design 12 metrics from three perspectives that are response metrics, global metrics, and node metrics, described as follows:
  • $\bullet$

### v3

@@
Abstract
Cryptocurrencies are widely used, yet current methods for analyzing transactions often rely on opaque, black-box models. While these models may achieve high performance, their outputs are usually difficult to interpret and adapt, making it challenging to capture nuanced behavioral patterns. Large language models (LLMs) have the potential to address these gaps, but their capabilities in this area remain largely unexplored, particularly in cybercrime detection. In this paper, we test this hypothesis by applying LLMs to real-world cryptocurrency transaction graphs, with a focus on Bitcoin, one of the most studied and widely adopted blockchain networks. We introduce a three-tiered framework to assess LLM capabilities: foundational metrics, characteristic overview, and contextual interpretation. This includes a new, human-readable graph representation format, LLM4TG, and a connectivity-enhanced transaction graph sampling algorithm, CETraS. Together, they significantly reduce token requirements, transforming the analysis of multiple moderately large-scale transaction graphs with LLMs from nearly impossible to feasible under strict token limits. Experimental results demonstrate that LLMs have outstanding performance on foundational metrics and characteristic overview, where the accuracy of recognizing most basic information at the node level exceeds 98.50% and the proportion of obtaining meaningful characteristics reaches 95.00%. Regarding contextual interpretation, LLMs also demonstrate strong performance in classification tasks, even with very limited labeled data, where top-3 accuracy reaches 72.43% with explanations. While the explanations are not always fully accurate, they highlight the strong potential of LLMs in this domain. At the same time, several limitations persist, which we discuss along with directions for future research.
Index Terms:
LLMs, Transaction Graph, Cybercrime Detection, Graph Representation, Cryptocurrency, Blockchain.
I Introduction
@@
Large language models (LLMs) [1] have significantly boosted the productivity of daily life and have a huge impact on the research community, such as in natural language processing (NLP) [2, 3], computer vision (CV) [4, 5], and application research [6, 7, 8]. The applications of LLMs also extend beyond traditional domains, influencing areas with social and economic implications.
@@
One such area is the cryptocurrency ecosystem. Its growing adoption in finance, retail, and entertainment has led to a surge in transaction volumes. However, the expansion also exposes the ecosystem to risks, such as scams and money laundering, enabled by its decentralized and pseudoanonymous nature. Current analysis methods rely on black-box models and struggle with interpretability of results and adaptability. In this context, applying LLMs to analyze cryptocurrency transactions offers a promising approach to bridging these gaps. By leveraging their capacity to interpret complex patterns and behaviors, LLMs can help identify illicit activities and enhance cybercrime detection efforts.
@@
Although LLMs trained on massive datasets excel in NLP tasks, their application to graph analysis presents challenges due to structural differences between graph and text data. Recent studies investigated the possibilities of LLMs for handling graph data-related tasks, concluding affirmatively that they are capable of completing specific tasks with acceptable performance on graphs such as small graphs, citation graphs, or knowledge graphs (KGs) [9, 10, 11, 12]. Nevertheless, measuring LLMs’ capability to understand and analyze cryptocurrency transaction graphs remains impractical. They contain different information compared with the other graph types such as KGs. Taking the Bitcoin transaction network as an example, the node represents the Bitcoin address or transaction, while the edge indicates the token flows among the address nodes and the transaction nodes [13].
@@
In addition, due to the input token limit of LLMs, how to efficiently feed larger graph data into LLMs to gain more information for potentially improving the quality of generated answers to various questions relevant to Bitcoin transaction graphs (e.g., address type prediction) continues to be an open question. To bridge the gaps in applying LLMs to transaction graph analysis, we study Bitcoin networks and address three research questions (RQs):
@@
Contributions. To the best of our knowledge, this is the first systematic study of LLMs’ capabilities in analyzing real-world cryptocurrency transaction graphs. We make the following contributions:
@@
  • We present a layered framework with three levels of understanding for measuring LLMs’ ability to analyze transaction graphs in cryptocurrency networks.
@@
  • We propose a text-based graph representation format, denoted LLM4TG. It reduces redundant data and provides a human-readable syntax that naturally supports processing by LLMs.
@@
  • We design a Connectivity-Enhanced Transaction Graph Sampling algorithm, CETraS, for graph summarization, removing less important nodes in moderately large-scale transaction graphs while enhancing critical connections.
  • We conduct both quantitative and qualitative evaluations of five representative models from three major LLM families (GPT, DeepSeek, and LLaMA) across multiple tasks, revealing their strengths and limitations in analyzing structural and behavioral patterns.
In the next section, we will review the extant literature.
II Related Work
II-A Cryptocurrency Transaction Analysis
Empirical Analysis. Empirical analysis plays an important role in understanding the dynamics of the cryptocurrency ecosystem [14, 15, 16] and the behaviors of addresses/entities [17, 18]. For instance, Tovanich et al. [19] and Hou et al. [20] reveal that factors such as payout schemes and pool fees influence miners’ behaviors in Bitcoin mining pools, and then impact the overall system performance.
Complex Network Analysis. Complex networks use graph theory, centrality measures, and network topology analysis to examine patterns within networks. This method explores the structures of cryptocurrency transaction networks from a macro perspective, highlighting the interconnections between nodes and the network’s overall structure. Nerurkar et al. [26] and Serena et al. [27] drew parallels between cryptocurrency systems and other complex systems and identified characteristics such as small-world property, indicating most nodes in the graph are not neighbors but most of them can be reached by every other within a few hops.
Moreover, Tao et al. [28] employed an innovative random walk with a flying-back sampling method on Bitcoin transaction graphs, uncovering phenomena such as the non-rich-club effect, i.e., that high-degree nodes are not more interconnected among themselves than with lower-degree nodes. Guo et al. [29] analyzed Ethereum transaction graphs, revealing heavy-tailed property in transaction networks, i.e., the majority of nodes have a relatively low degree while a small number of nodes have a very high degree. To mitigate the issue of money laundering on blockchain networks, DenseFlow framework, proposed by Lin et al. [30], uses dense subgraphs and the maximum flow algorithm to trace laundering activities. This approach improves precision compared to existing methods on Ethereum, demonstrating the effectiveness of network analysis in combating money laundering.
Machine Learning Analysis. Machine learning is used to achieve node- or graph-level classification and prediction tasks for concrete addresses. Chaudhari et al. [31] studied utilizing temporal features to detect Bitcoin address behavioral changes and identify money laundering activities. The proposed approaches based on the decision tree (DT) by Rathore et al. [32] show high accuracy rates in detecting illicit activities and phishing scams in cryptocurrencies. Wahrstatter et al. [33] also contributed by enhancing the detection of criminal activities in Bitcoin transactions using unsupervised learning.
Additionally, various machine learning methods, such as random forest (RF), multilayer perceptron (MLP), and graph neural network (GNN), are applied to benchmark cryptocurrency datasets, including the Elliptic Data Set by Weber et al. [34] and datasets by Xiang et al. [13, 35]. Besides, Gai et al. [36] proposed a transformer-based anomaly detection model BlockGPT for the Ethereum network that demonstrated acceptable utility.
II-B LLMs in Graph Analysis and Current Gaps
LLMs in Graph Analysis. Several studies introduced LLMs to graph analysis, using LLM as a classifier and GNN enhancement [37]. Wang et al. [11, 38] evaluated the basic capabilities of LLMs in natural language graph problem-solving. Both studies showed limitations of LLMs, particularly in solving complex graph structures and tasks. Complementing these insights, Tang et al. [39] and Hu et al. [10] assessed the performance of LLMs in graph data analysis and prediction, compared them with specialized GNNs. Likewise, Sui et al. [40] and Jiang et al. [41] explored the effectiveness of LLMs including GPT-3.5 and GPT-4 in processing structured data, such as tables and various structured data types, introducing innovative prompting methods for performance enhancement.
Current Gaps. We conclude three primary research limitations according to the abovementioned work:
  • Existing research mainly focuses on knowledge graphs and randomly generated graphs [42, 10, liu2023evaluating, 11]. However, how to measure the LLMs’ ability to understand and analyze real-world cryptocurrency transaction graphs is unresolved.
Figure 1: LLM Evaluation Framework for Bitcoin Transaction
III Preliminaries
@@
Transaction Graphs for Cryptocurrencies. Cryptocurrency transaction graphs [17, 34, 29, 38] represent the flow of digital currency between entities on blockchain networks such as Bitcoin. Each node in the graph typically represents a transaction or account/address, and edges represent the transfer of cryptocurrency between these nodes. These graphs are crucial for analyzing the behavior of users, identifying patterns such as fraud or money laundering, and understanding the overall dynamics of the cryptocurrency market.
@@
LLMs for Structured Data Analysis. LLMs [1] like OpenAI’s Generative Pre-trained Transformer (GPT) [46] series are advanced models trained on vast amounts of text data. They excel in generating coherent and contextually relevant text based on the input they receive and can perform a variety of tasks without specific task-oriented training. While primarily designed for natural language processing, LLMs can be adapted to non-textual tasks such as analyzing structured data, including graphs [47, 37]. By converting data into a format that mimics natural language or structured prompts, researchers can leverage LLMs’ powerful generative and interpretative capabilities to perform complex analyses like those required for understanding cryptocurrency transaction graphs.
@@
Token Limitations in LLMs. The term token indicates the smallest unit of text processed by LLMs. It can vary from individual characters, such as letters and punctuation, to more complicated units, such as words or subwords. LLMs analyze and generate text by processing token sequences. Each LLM has a maximum token limit per input, which poses a significant challenge when working with extensive data sets, such as large transaction graphs. The token limit in LLMs [46] constrains data analysis in a single query and poses significant challenges for analyzing large and complex transaction graphs, potentially leading to loss of context or omission of critical structures. Although strategies such as compression and sampling have been explored, they often compromise structural integrity or contextual completeness, especially in moderately large-scale graphs.
IV Methodology
IV-A Framework Overview
@@
We present our framework in Fig. 1. The first step is to construct the Bitcoin transaction graph using the historical on-chain data. The second step is to extract corresponding subgraphs (or the relevant subgraphs sampled by CETraS, and graph features to Bitcoin addresses with labels. Finally, raw graphs formatted as LLM4TG and graph features are input to LLMs with proper prompts according to the tasks on different understanding levels.
IV-B LLM4TG
@@
We introduce LLM4TG, a new graph representation format designed to optimize the analysis of transaction graphs using LLMs. This format is text-based and human-readable, minimizing syntactic noise/redundancy while reducing token usage and preserving data integrity.
@@
We denote T as transaction and A as address. <NodeID> represents the node’s ID, <Number> represents an integer. <Float> represents a real number. The syntax is displayed in Listing 1.
This format provides three key advantages for representing and analyzing transaction graphs: 1) It organizes nodes into type-specific layers, mirroring the structure of original transaction graphs; 2) It efficiently utilizes the limited token budget of LLMs by allowing more data to be encoded; 3) It improves the interpretability of graph data for LLMs by organizing node attributes into closely associated key-value pairs.
To further demonstrate the effectiveness of LLM4TG, we compared its token consumption with other formats for the same graphs, as shown in Fig. 7. This comparison reveals that LLM4TG experiences a more gradual increase in token usage and consistently stays within the GPT-4/4o token limit across various graph sizes, making it a more efficient format, especially for larger graphs. Further details and discussion are presented in Section VI.
@@
IV-C CETraS
@@
In CETraS, nodes with lower importance are prioritized for elimination.to generate a subset of the nodes being preserved. The size of this retained subset is determined by a parameter that is flexible for specific demands. To maintain connectivity, paths connecting retained nodes are also preserved. Unlike other state-of-the-art graph summarization methods [48, 49, 50, 51] focusing on keeping accuracy for structure-relevant queries or computations on large-scale graphs (billion node-level), CETraS concentrates on accurately conveying transaction-relevant information to LLMs for moderately large-scale graphs that typically contain thousands of nodes.
@@ 3 $I_{node}\leftarrow\left\{v:I_{node}[v]\text{ for each }v\text{ in }V(G)\right\}$;
4 $P_{node}\leftarrow\frac{1}{I_{node}}$;
5 $P_{n_{0}}\leftarrow 0$;
6 $P_{sum}\leftarrow\sum P_{node}$;
7 foreach node $n$ $\in$ $G$ do
8    $P_{n}\leftarrow\frac{P_{n}}{P_{sum}}$;
9
10   end foreach
11 $G_{subset}\leftarrow$ Sample from $V(G)$ with $P_{node}$ until $|G_{subset}|\geq\min\{N_{target},|V(G)|\}$ ;
12
@@ 13 $G_{sampled}\leftarrow$ Initialize an empty graph;
14 foreach node $n$ $\in$ $G_{subset}$ do
@@ 17         Add node $m$ to $G_{sampled}$;
20    foreach edge $e$ $\in$ $p$ do
21         Add edge $e$ to $G_{sampled}$;
22
23      end foreach
24
@@ 21         Add edge $e$ to $G_{sampled}$;
V Evaluation and Analysis
V-A Settings
Dataset. We use two datasets for experiments, both constructed from the whole Bitcoin transaction graph, which spans a 22-month period (12 July 2019 – 26 May 2021). Specifically, BASD [35] includes eight types of subgraphs, each starting from a labeled address (denoted as $n_{0}$) and extending up to five hops with at most 3,000 nodes. It is one of the few datasets that provides structured graph-level representations of Bitcoin addresses. Meanwhile, BABD [13] contains labeled Bitcoin addresses, each associated with 148 engineered features. It is one of the largest publicly available datasets for Bitcoin address behavior analysis.
LLM Selection. We select five LLMs for evaluation: GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-turbo), GPT-4o (gpt-4o), DeepSeek (deepseekv3), and LLaMA (llama3.3-70B). These models are chosen for their strong overall performance and representativeness in specific tasks. Due to differences in their input length capacities (e.g., up to 8,192 tokens for llama3.3-70B), not all models are evaluated on every task. All models are accessed via application programming interfaces (APIs), which support prompt-based querying and response generation.
@@ V-A Settings
Graph Formatting and Model Allocation. We format raw transaction graphs as LLM4TG in the experiments. We use CETraS in level 2 and level 3 due to the token limit in few-shot prompts. We apply GPT-3.5 only in level 3 feature-based classification, given its particularly tight token limits (see Fig.7). Experiments are conducted on BASD dataset subsets and its corresponding BABD addresses.
Prompt Engineering and Design. We employ few-shot prompting [52] to interact with LLMs, providing only a small number of examples to guide task execution. To further align prompts with the model’s capabilities and preferences, we use an LLM-feedback generation strategy, wherein the model is queried with task-specific requirements and subsequently designs the corresponding system prompts. All code, prompts, and results are publicly available at our GitHub. [^1] https://github.com/yuchen-lei/llm4tg
Experimental Setup. All experiments are conducted on a workstation equipped with an AMD Ryzen Threadripper PRO 5995WX and 256 GB of RAM, running Ubuntu 22.04.
V-B Level 1 - Foundational Metrics
We evaluate LLMs on sampled transaction graphs with selected nodes across layers. The evaluation adopts twelve metrics covering three perspectives: response metrics, global metrics, and node metrics: